OpenStudy (anonymous):

Please help I will give Medal... Evaluate each of the following: sin^120° + cos^2 120°

OpenStudy (diamondboy):

\[\sin^2120 + \cos^2120\]

OpenStudy (diamondboy):

is dat ur question

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

Rule: \[\Large \sin^2(x) + \cos^2(x) = 1\] is a true equation for any real number x value. This is known as an identity. Specifically the pythagorean identity http://www.regentsprep.org/regents/math/algtrig/att9/pythagoreanid.htm

OpenStudy (diamondboy):

d answer is 1

OpenStudy (anonymous):

Really I Did it like this but I didnt know how to contrinue I did like this( √3/2)^2 + (- 1/2)^2= 3/2 + 1/4

jimthompson5910 (jim_thompson5910):

( √3/2)^2 will turn into 3/4 you square both the √3 and the 2 down below

jimthompson5910 (jim_thompson5910):

3/4 + 1/4 = 4/4 = 1

OpenStudy (anonymous):

Ok get it :) thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Can u help me with another one really similar?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure